Shadee Corp. expects to sell 510 sun visors in May and 420 in June. Each visor sells for $21. Shadee’s beginning and ending finished goods inventories for May are 80 and 45 units, respectively. Ending finished goods inventory for June will be 70 units.

1. Determine Shandees budgeted Total Sales for May and June

2. Determine Shandees Budgeted Production for May and June
Answer #1

Budgeted total sales for

May = 510 * 21 = 10,710

June = 420 * 21 = 8,820

Budgeted production = Budgeted sales + Ending inventory - Beginning inventory

May = 510 + 45 - 80 = 475 units

​​​​​​June = 420 + 70 - 45 = 445 units

Comment if you face any issues